OpenStudy (anonymous):

What is 3.2 × 10−4 m/h in centimeters per second? A. 8.889 × 10^−6 cm/s B. 8.889 × 10^−5 cm/s C. 1.152 × 10^−5 cm/s D. 1.152 × 10^−7 cm/s

OpenStudy (wolf1728):

3.2 x 10^-4 m/h /3,600 = 8.888E-008 m/sec 8.888E-008 m/sec x 100 cm/m = 8.888-006 cm/sec So it seems the answer is A